# Towards Tight Convex Relaxations for Contact-Rich Manipulation

### v1

@@ Abstract
We present a method for global motion planning of robotic systems that interact with the environment through contacts. Our method directly handles the hybrid nature of such tasks using tools from convex optimization. We formulate the motion-planning problem as a shortest-path problem in a graph of convex sets, where a path in the graph corresponds to a contact sequence and a convex set models the quasi-static dynamics within a fixed contact mode. For each contact mode, we use semidefinite programming to relax the nonconvex dynamics that results from the simultaneous optimization of the object’s pose, contact locations, and contact forces. The result is a tight convex relaxation of the overall planning problem, that can be efficiently solved and quickly rounded to find a feasible contact-rich trajectory. As a first application of this technique, we focus on the task of planar pushing. Exhaustive experiments show that our convex-optimization method generates plans that are consistently within a small percentage of the global optimum. We demonstrate the quality of these plans on a real robotic system.
@@ I Introduction
Optimal planning and control through contact is an important and challenging problem in robotics, which includes tasks like robot locomotion and dexterous manipulation. It generally involves both a hybrid and underactuated dynamical system, making planning and control difficult. A variety of methods have been proposed to tackle this problem. Many of them artificially decouple the discrete and continuous components of the problem, yielding suboptimal trajectories that do not take full advantage of the rich contact dynamics. Approaches that blend discrete and continuous often do so locally (around a given trajectory) and are unable to reason about the global problem; or rely on expensive global optimization algorithms that scale poorly.
@@
In this work, we introduce a method that naturally blends the discrete logic and the continuous dynamics of planning through contact into a convex optimization problem. Specifically, we formulate the planning problem as a Shortest Path Problem (SPP)in a Graph of Convex Sets (GCS), a class of mixed discrete/continuous optimization problems that can be effectively solved to global optimality [1]. Paths in this graph correspond to different contact sequences, and the convex sets model the quasi-static contact dynamics within a fixed contact mode. These quasi-static dynamics are bilinear (therefore nonconvex), since we simultaneously optimize over the object pose, contact locations, and contact forces. Our method approximates these bilinearities using a tight Semidefinite Programming (SDP)relaxation for each contact mode. The methods from [1] are then leveraged to produce a tight convex relaxation of the global planning-through-contact problem. This relaxation can be quickly solved and rounded to a feasible contact-rich trajectory. By comparing the cost of the rounded solution and the convex relaxation, our motion planner also provides us with tight optimality bounds for the trajectories that it designs.
As a first application of our method, this work explores the task of planar pushing, first studied by Mason in [2]. Planar pushing has applications that span from warehouse automation to service robotics, and although it is among the simplest examples of non-prehensile manipulation, current state-of-the-art approaches are still unable to solve this problem reliably to global optimality. The method proposed in this paper can simultaneously reason about the high-level discrete mode switches and the low-level continuous dynamics of planar pushing. We evaluate our motion planner through thorough numerical experiments, which show that the trajectories we generate typically have a very small optimality gap ($10\%$ on average). We also demonstrate our approach on the real robotic system shown in in figure 1. We emphasize that, although we study the tightness of our relaxations on the particular domain of planar pushing, the technique we introduce generalizes naturally to more complex multi-contact problems.
@@ II Related Work
The learning methods that have been proven most effective for contact-rich tasks are Reinforcement Learning (RL)and Behavior Cloning (BC). RL has proven particularly effective due to its ability to bypass the low-level combinatorial contact decisions by using stochastic approximation of the system’s dynamics [3, 4, 5]. In particular, RL has shown impressive robustness for real-world quadruped locomotion [6], and has been successfully applied to manipulation tasks in simulation [7, 8], real-world robotic manipulation [9, 10], and more recently on real-world dexterous hands [11, 12, 13, 14]. Similarly, BC methods have been successfully applied to many real-world manipulation tasks [15, 16, 17, 18], including planar pushing [15, 19]. Although RL and BC are applicable to a variety of systems, RL methods tend to require expensive training and task-specific user-shaped reward functions, while BC requires costly expert demonstrations that are typically generated by human teleoperation. The method we propose here can potentially be used to automatically generate large sets of demonstrations of super-human quality.
@@
Global optimization methods for planning through contact typically use Mixed-Integer Programming (MIP)to jointly optimize over discrete contact modes and continuous robot motions. A common approach is to produce a Piecewise Affine (PWA)approximation of the nonlinear dynamics, casting the problem as an MIP [38]. This and similar approaches have been applied to push-recovery [39], locomotion [40], and planar pushing [41, 42]. Recently, efficient task-specific PWA approximation have also been obtained by pruning large neural networks [43]. The main limitation of these approaches is the MIP runtime, which can grow exponential with the number of integer variables. In addition, PWA models are also unable to accurately capture the smooth contact dynamics within a fixed contact mode. At the cost of a further slowing down computations, the latter issue can be tackled by introducing additional binary variables (see, e.g., [44]). Alternatively, by relying on expensive mixed-integer nonconvex solvers [45] or tailored convex relaxations [46].
Our method is similar to the ones in the last category, but is substantially more efficient. We depart from prior work by leveraging a tight convex relaxation for the bilinearities in each contact mode, while also leveraging the methods from [1] to produce a tight convex relaxation for the global planning-through-contact problem. The tightness of both relaxations enables us to obtain near-globally optimal solutions by solving the entire problem as a single SDP, followed by a quick rounding step. This is the first work to our knowledge that is able to effectively solve the global planning-through-contact problem as a single convex program.
@@ III Problem Statement
As a first application of our method, we explore planar pushing, a non-prehensile manipulation task where the robot uses a cylindrical finger to manipulate the motion of an object resting on a flat surface. We shall refer to the manipulated object as the slider, the robot finger as the pusher, and to the overall system as the slider-pusher system. We assume that the slider and the pusher are rigid bodies, and restrict the problem to planar motion, such that the slider is always making contact with the underlying surface. We also assume that the slider is a prism-shaped object with known geometry, which can be non-convex. See Figure 1 for a picture of the the real-world problem setup.
As the slider is pushed, it is subject to a contact wrench from the pusher and a friction wrench from the underlying surface (we assume a uniform pressure distribution between the slider and the surface). We assume quasi-static dynamics (i.e., no work is done by impacts) and that the velocities are sufficiently low to make inertial forces negligible. Under this assumption, for the slider to move, the contact wrench applied by the pusher must balance the friction wrench from the underlying surface. We assume isotropic Coulomb friction, i.e., that the coefficient of friction is constant, and the friction force at every contact point must have a constant magnitude and oppose the direction of motion during sliding. These assumptions are standard in the non-prehensile literature [47, 48, 49].
@@ IV High-Level Approach
The first step in formulating our motion planning method is to consider the dynamics and kinematics in a fixed contact mode. By formulating the manipulation problem in the task space, we are able to describe the system with only quadratic and linear constraints. More specifically, quadratic equality constraints arise from considering elements of $\mathrm{SO}(2)$, rotations of velocities and forces to the world frame, and the contact torque as the cross-product between the contact arm and force. We thus formulate the motion planning problem within a specific contact mode as a non-convex quadratically constrained, quadratic program (QCQP), which we approximate as a convex program using a semidefinite relaxation.
@@ V-A Shortest paths in graph of convex sets
In this section, the formulation for the SPP in a GCS is briefly reviewed. For further details, the reader is referred to [1]. We define a Graph of Convex Sets (GCS) as a directed graph $G=(\mathcal{V},\mathcal{E})$ with vertex set $\mathcal{V}$ and edge set $\mathcal{E}\subset\mathcal{V}^{2}$, where each vertex $v\in\mathcal{V}$ is paired with a compact convex set $\mathcal{X}_{v}$ and a point $x_{v}$ contained in the set. Additionally, each vertex $v$ is associated with a vertex cost $l_{v}(x_{v})$, a nonnegative convex function of the point $x_{v}\in\mathcal{X}_{v}$. Finally, the edges of the graph are associated to convex constraints of the form $(x_{u},x_{v})\in\mathcal{X}_{e}$ which couple the vertices in an edge $e$.
@@
Problem (V-A) can be solved exactly as a Mixed-Integer Convex Program (MICP). Importantly, the problem formulation has a very tight convex relaxation, and can in many instances be solved to global optimality by solving the convex relaxation and performing a cheap rounding step on the integer variables, as shown in [50]. In this work, we treat the GCS framework as a modelling language that allows us to formulate and efficiently solve the problem formulated in (V-A). Additionally, the GCS framework naturally gives us an upper bound on the optimality gap to a solution: Let $C_{\text{relax}}\leq C_{\text{opt}}\leq C_{\text{round}}$ be the costs of the relaxation, the MICP, and the rounded solution, respectively. The optimality gap $\delta_{\text{opt}}$ can the be overestimated as
@@ V-B Semidefinite relaxation of quadratically constrained quadratic programs
In this section we review non-convex QCQPs and their semidefinite relaxation [51], and refer the reader to [52, 53, 54] for a more detailed discussion. Let $y\in\mathbb{R}^{n}$ and let $x\coloneqq\begin{bmatrix}1&y^{\intercal}\end{bmatrix}^{\intercal}$. A non-convex QCQP in homogeneous form is the optimization program
@@
where $A\in\mathbb{R}^{m\times(n+1)}$, and $Q_{i}\in\mathbb{R}^{(n+1)\times(n+1)},\,i=0,\ldots,l$. Note that positive-semidefiniteness is not required for $Q_{i}$, hence (V-B) can be non-convex.
A lower bound to (V-B) can be obtained by solving the semidefinite program [51, 53]
@@
$X$ $=\begin{bmatrix}1&y^{\intercal}\\ y&Y\end{bmatrix}\succeq 0$ (4g)
@@
Recalling that if the PSD matrix $X$ is rank one, then $X=xx^{T}$, and using the simple identity $x^{\intercal}Q_{i}x=\operatorname{tr}{(Q_{i}xx^{\intercal})}$, it can be seen that any feasible, rank-one solution to (V-B) provides a feasible solution (V-B). Therefore (V-B) is indeed a relaxation. Moreover, if $X^{*}$ is a rank-one minimizer of (V-B), then $x^{*}$ satisfying $X^{*}=x^{*}(x^{*})^{\intercal}$ is the global optimum of (V-B). This motivates the following scheme for obtaining near-optimal solutions to (V-B). After solving (V-B) to obtain an $X^{*}$, we take $\hat{x}=\begin{bmatrix}1&\hat{y}^{\intercal}\end{bmatrix}^{\intercal}$ to be the first column of $X^{*}$. We then use $\hat{x}$ as an initial guess for solving (V-B) using a local non-convex solver. If a feasible solution is found, then we have the bound $C_{\text{relax}}\leq C_{\text{opt}}\leq C_{\text{round}}$ where $C_{\text{relax}}$, $C_{\text{opt}}$, and $C_{\text{round}}$ are the optimal costs of the semidefinite relaxation, the non-convex QCQP, and the rounded solution, respectively. While this strategy is not guaranteed to find a feasible solution to (V-B), our results in Section VIII demonstrate that in the case of planar pushing, this strategy almost always finds both feasible and near-optimal solutions.
This semidefinite relaxation and related techniques are commonly used in combinatorial optimization (see, e.g., [55]). They have been used in robotics for localization and pose estimation [56, 57, 58] and motion planning around obstacles [59].
@@ VI The Slider-Pusher System
We use the monogram notation from [60, §3.1] for reference frames. Denote the position of frame $B$ relative to $A$, measured in frame $C$ by $\prescript{A}{}{p}^{B}_{C}$ (when $C=A$, we write simply $\prescript{A}{}{p}^{B}$). We similarly write $\prescript{A}{}{v}^{B}_{C}$ (or $\prescript{A}{}{v}^{B}$) for translational velocities, and $\prescript{A}{}{V}^{B}_{C}$ (or $\prescript{A}{}{V}^{B}$) for spatial velocities. The rotation that transforms a vector from frame $B$ to frame $A$ is denoted by $\prescript{A}{}{R}^{B}$. To denote a Cartesian force applied at a point $p$ measured in frame $A$ we will write $f^{p}_{A}$, and to denote the spatial force we write $F^{p}_{A}$. We will drop the frames when they are clear from the context.
@@ VI-A Kinematics
Consider the slider-pusher system in Figure 1 (a). Let $W$ be the world frame, and attach the reference frames $S$ and $P$ at the COM of the slider and pusher, respectively. Let the pose of the slider in the world frame be given by $(\prescript{W}{}{p}^{S},\prescript{W}{}{\theta}^{S})\in\mathrm{SE}(2)$. Let the (planar) spatial velocity of the slider be $\prescript{W}{}{V}^{S}=(\prescript{W}{}{v}^{S},\prescript{W}{}{\omega}^{S})\in\mathrm{se}(2)$ (the tangent space associated with $\mathrm{SE}(2)$), where $\prescript{W}{}{v}^{S}=(d/dt)\prescript{W}{}{p}^{S}$ and $\prescript{W}{}{\omega}^{S}=(d/dt)\prescript{W}{}{\theta}^{S}$. We represent the position of the pusher in the frame of the slider by $\prescript{S}{}{p}^{P}\in\mathbb{R}^{2}$, and denote its velocity by $\prescript{S}{}{v}^{P}=(d/dt)\prescript{S}{}{p}^{P}$. We do not assign an orientation to the pusher, as it is assumed to be circular and hence invariant to rotation. When there is contact between the slider and the pusher, we will denote the position of the contact point in the slider frame by $\prescript{S}{}{p}^{c}\in\mathbb{R}^{2}$ and the velocity by $\prescript{S}{}{v}^{c}=(d/dt)\prescript{S}{}{p}^{c}$ Going forward we will use these definitions, but will often drop the frames to keep notation light. We use $(p^{S},\theta)$ and $V=(v^{S},\omega^{S})$ to denote the slider pose and spatial velocity in the world frame, and $p^{P}$, $v^{P}$, $p^{c}$, and $v^{c}$ to denote the pusher position, translational pusher velocity, contact point position, and contact point velocity in the frame of the slider.

### v2

@@ Abstract
We present a novel method for global motion planning of robotic systems that interact with the environment through contacts. Our method directly handles the hybrid nature of such tasks using tools from convex optimization. We formulate the motion-planning problem as a shortest-path problem in a graph of convex sets, where a path in the graph corresponds to a contact sequence and a convex set models the quasi-static dynamics within a fixed contact mode. For each contact mode, we use semidefinite programming to relax the nonconvex dynamics that results from the simultaneous optimization of the object’s pose, contact locations, and contact forces. The result is a tight convex relaxation of the overall planning problem, that can be efficiently solved and quickly rounded to find a feasible contact-rich trajectory. As an initial application for evaluating our method, we apply it on the task of planar pushing. Exhaustive experiments show that our convex-optimization method generates plans that are consistently within a small percentage of the global optimum, without relying on an initial guess, and that our method succeeds in finding trajectories where a state-of-the-art baseline for contact-rich planning usually fails. We demonstrate the quality of these plans on a real robotic system.
@@ I Introduction
Optimal planning and control through contact is an important and challenging problem in robotics, which includes tasks like robot locomotion and dexterous manipulation. It generally involves both a hybrid and underactuated dynamical system, making planning and control difficult. A variety of methods have been proposed to tackle this problem. Many of them artificially decouple the discrete and the continuous components of the problem, yielding suboptimal trajectories that do not take full advantage of the rich contact dynamics. Approaches that blend the discrete and continuous components often do so locally (around a given trajectory) and are unable to reason about the global problem; or rely on expensive global optimization algorithms that scale poorly.
@@
In this work, we introduce a method that naturally blends the discrete logic and the continuous dynamics of planning through contact into a convex optimization problem. Specifically, we formulate the planning problem as a Shortest Path Problem (SPP)in a Graph of Convex Sets (GCS), a class of mixed discrete/continuous optimization problems that can be effectively solved to global optimality [1]. Paths in this graph correspond to different contact sequences, and the convex sets model the quasi-static contact dynamics within a fixed contact mode. These quasi-static dynamics are bilinear (therefore nonconvex), since we simultaneously optimize over the object pose, contact locations, and contact forces. Our method approximates these bilinearities using a tight Semidefinite Programming (SDP)relaxation for each contact mode. The methods from [1] are then leveraged to produce a tight convex relaxation of the global planning-through-contact problem. This relaxation can be quickly solved and rounded to obtain a feasible contact-rich trajectory. By comparing the cost of the rounded solution and the convex relaxation, our motion planner also provides us with tight optimality bounds for the trajectories that it designs.
As a first application for evaluating our method, this work explores the task of planar pushing, first studied by Mason in [2]. Planar pushing has applications that span from warehouse automation to service robotics, and although it is among the simplest examples of non-prehensile manipulation, current state-of-the-art approaches are still unable to solve this problem reliably to global optimality. The method proposed in this paper can simultaneously reason about the high-level discrete mode switches and the low-level continuous dynamics of planar pushing in a global fashion. We evaluate our motion planner through thorough numerical experiments, which show that the trajectories we generate typically have a very small optimality gap ($10\%$ on average). We also demonstrate our approach on the real robotic system shown in in Figure 1. We emphasize that, although we study the tightness of our relaxations on the particular domain of planar pushing, the technique we introduce generalizes naturally to more complex multi-contact problems.
@@ II Related Work
The learning methods that have been proven most effective for contact-rich tasks are Reinforcement Learning (RL)and Behavior Cloning (BC). RL has proven particularly effective due to its ability to bypass the low-level combinatorial contact decisions by using stochastic approximation of the system’s dynamics [3, 4, 5]. In particular, RL has shown impressive robustness for real-world quadruped locomotion [6], and has been successfully applied to manipulation tasks in simulation [7, 8], real-world robotic manipulation [9, 10], and more recently on real-world dexterous hands [11, 12, 13, 14]. Similarly, BC methods have been successfully applied to many real-world manipulation tasks [15, 16, 17, 18], including planar pushing [15, 19]. Although RL and BC are applicable to a variety of systems, RL methods tend to require expensive training and task-specific user-shaped reward functions, while BC requires costly expert demonstrations that are typically generated by human teleoperation. The method we propose in this paper can potentially be used to automatically generate large sets of expert demonstrations of very high quality for a variety of different tasks.
@@
Global optimization methods for planning through contact typically use Mixed-Integer Programming (MIP)to jointly optimize over discrete contact modes and continuous robot motions. A common approach is to produce a Piecewise Affine (PWA)approximation of the nonlinear dynamics, casting the problem as an MIP [38]. This and similar approaches have been applied to push-recovery [39], locomotion [40], and feedback control for planar pushing [41, 42]. Recently, efficient task-specific PWA approximation have also been obtained by pruning large neural networks [43]. The main limitation of these approaches is the MIP runtime, which can grow exponentially with the number of integer variables. In addition, PWA models are also unable to accurately capture the smooth contact dynamics within a fixed contact mode. At the cost of a further slowing down computations, the latter issue can be tackled by introducing additional binary variables and approximating the nonconvexities with piecewise linear functions (see, e.g., [44] or [45]), possibly in combination with tailored convex relaxations [46]. Alternatively, general Mixed-Integer Nonconvex Programming (MINCP)approaches have been used in robotics for momentum-based footstep planning [47], but their applicability is limited to very small problem instances.
Our method is similar to the ones in the last category, but is substantially more efficient. Rather than resorting to MINCP or piecewise linear approximations combined with mixed-integer programming, we depart from prior work by leveraging a tight convex relaxation for the bilinearities in each contact mode while also leveraging the methods from [1] to produce a tight convex relaxation for the global planning-through-contact problem. The tightness of both relaxations enables us to obtain near-globally optimal solutions by solving the entire problem as a single SDP, followed by a quick rounding step. This is the first work to our knowledge that is able to effectively solve the global planning-through-contact problem as a single convex program.
@@ III Problem Statement
As a first application of our method, we explore planar pushing, a non-prehensile manipulation task where the robot uses a cylindrical finger to manipulate the motion of an object resting on a flat surface. We shall refer to the manipulated object as the slider, the robot finger as the pusher, and to the overall system as the slider-pusher system. We assume that the slider and the pusher are rigid bodies, and restrict the problem to planar motion, such that the slider is always making contact with the underlying surface. We also assume that the slider is a prism-shaped object with known geometry, which can be nonconvex, and a known (but arbitrary) COM. See Figure 1 for a picture of the the real-world problem setup.
As the slider is pushed, it is subject to a contact wrench from the pusher and a friction wrench from the underlying surface (we assume a uniform pressure distribution between the slider and the surface). We assume quasi-static dynamics (i.e., no work is done by impacts) and that the velocities are sufficiently low to make inertial forces negligible. Under this assumption, for the slider to move, the contact wrench applied by the pusher must balance the friction wrench from the underlying surface. We assume isotropic Coulomb friction, i.e., that the coefficient of friction is constant, and the friction force at every contact point must have a constant magnitude and oppose the direction of motion during sliding. Although these assumptions are common in the literature on non-prehensile modelling [48, 49] and planning [50, 51], we note that a more complex friction model could in principle be incorporated directly into our framework or as a post-processing step.
@@ IV High-Level Approach
The first step in formulating our motion planning method is to consider the dynamics and kinematics in a fixed contact mode. By formulating the manipulation problem in the task space, we are able to describe the system with only quadratic and linear constraints. More specifically, quadratic equality constraints arise from considering elements of $\mathrm{SO}(2)$, rotations of velocities and forces to the world frame, and the contact torque as the cross-product between the contact arm and force. We thus formulate the motion planning problem within a specific contact mode as a nonconvex Quadratically Constrained Quadratic Program (QCQP), which we approximate as a convex program using a semidefinite relaxation.
@@ V-A Shortest paths in graph of convex sets
In this section, we briefly review the formulation for the SPP in a GCS. For further details, the reader is referred to [1]. We define a Graph of Convex Sets (GCS) as a directed graph $G=(\mathcal{V},\mathcal{E})$ with vertex set $\mathcal{V}$ and edge set $\mathcal{E}\subset\mathcal{V}^{2}$, where each vertex $v\in\mathcal{V}$ is paired with a compact convex set $\mathcal{X}_{v}$ and a point $x_{v}$ contained in the set. Additionally, each vertex $v$ is associated with a vertex cost $l_{v}(x_{v})$, a nonnegative convex function of the point $x_{v}\in\mathcal{X}_{v}$. Finally, the edges of the graph are associated to convex constraints of the form $(x_{u},x_{v})\in\mathcal{X}_{e}$ which couple the vertices in an edge $e$.
@@
Problem (1) can be solved exactly as a Mixed-Integer Convex Program (MICP). Importantly, the problem formulation has a very tight convex relaxation, and can in many instances be solved to global optimality by solving the convex relaxation and performing a cheap rounding step on the integer variables, as shown in [52]. In this work, we treat the GCS framework as a modelling language that allows us to formulate and efficiently solve the problem presented in (1). Additionally, the GCS framework naturally gives us an upper bound on the optimality gap to a solution; Let $C_{\text{relax}}\leq C_{\text{opt}}\leq C_{\text{round}}$ be the costs of the relaxation, the MICP, and the rounded solution, respectively. The optimality gap $\delta_{\text{opt}}$ can then be overestimated as
@@ V-B Semidefinite relaxation of quadratically constrained quadratic programs
In this section we review nonconvex QCQPs and their semidefinite relaxation [53], and refer the reader to [54, 55, 56] for a more detailed discussion. This semidefinite relaxation and related techniques are commonly used in combinatorial optimization (see, e.g., [57]), and have been used in robotics for both localization and pose estimation [58, 59, 60], as well as motion planning around obstacles [61].
Let $y\in\mathbb{R}^{n}$ and let $x\coloneqq\begin{bmatrix}1&y^{\intercal}\end{bmatrix}^{\intercal}$. A nonconvex QCQP in homogeneous form is the optimization program
@@
where $A\in\mathbb{R}^{m\times(n+1)}$, and $Q_{i}\in\mathbb{R}^{(n+1)\times(n+1)},\,i=0,\ldots,l$. Note that positive-semidefiniteness is not required for $Q_{i}$, hence (3) can be nonconvex. A lower bound to (3) can be obtained by solving the semidefinite program [53, 55]:
@@
$X$ $=\begin{bmatrix}1&y^{\intercal}\\ y&Y\end{bmatrix}\succeq 0$ (4e)
@@
Recalling that if the PSD matrix $X$ is rank one, then $X=xx^{T}$, and using the simple identity $x^{\intercal}Q_{i}x=\operatorname{tr}{(Q_{i}xx^{\intercal})}$, it can be seen that any feasible, rank-one solution to (4) provides a feasible solution to (3). Therefore (4) is indeed a relaxation. Moreover, if $X^{*}$ is a rank-one minimizer of (4), then $x^{*}$ satisfying $X^{*}=x^{*}(x^{*})^{\intercal}$ is the global optimum of (3). This motivates the following scheme for obtaining near-optimal solutions to (3): After solving (4) to obtain an $X^{*}$, we take $\hat{x}=\begin{bmatrix}1&\hat{y}^{\intercal}\end{bmatrix}^{\intercal}$ to be the first column of $X^{*}$. We then use $\hat{x}$ as an initial guess for solving (3) using a local nonconvex solver. If a feasible solution is found, then we have the bound $C_{\text{relax}}\leq C_{\text{opt}}\leq C_{\text{round}}$ where $C_{\text{relax}}$, $C_{\text{opt}}$, and $C_{\text{round}}$ are the optimal costs of the semidefinite relaxation, the nonconvex QCQP, and the rounded solution, respectively. While this strategy is not guaranteed to find a feasible solution to (3), our results in Section VIII demonstrate that in the case of planar pushing, this strategy finds both feasible and near-optimal solutions for all the tested problem instances.
@@ VI The Slider-Pusher System
We use the monogram notation from [62, §3.1] for reference frames. Denote the position of the frame $B$ relative to $A$, measured in frame $C$ by $\prescript{A}{}{p}^{B}_{C}$ (when $C=A$, we write simply $\prescript{A}{}{p}^{B}$). We similarly write $\prescript{A}{}{v}^{B}_{C}$ (or $\prescript{A}{}{v}^{B}$) for translational velocities, and $\prescript{A}{}{V}^{B}_{C}$ (or $\prescript{A}{}{V}^{B}$) for spatial velocities. The rotation that transforms a vector from frame $B$ to frame $A$ is denoted by $\prescript{A}{}{R}^{B}$. To denote a Cartesian force applied at a point $p$ measured in frame $A$ we will write $f^{p}_{A}$, and to denote the spatial force we write $F^{p}_{A}$. We will drop the frames when they are clear from the context.
@@ VI-A Kinematics
Consider the slider-pusher system in Figure 2 (a). Let $W$ be the world frame, and attach the reference frames $S$ and $P$ at the COM of the slider and pusher, respectively. Let the pose of the slider in the world frame be given by $(\prescript{W}{}{p}^{S},\prescript{W}{}{\theta}^{S})\in\mathrm{SE}(2)$. Let the (planar) spatial velocity of the slider be $\prescript{W}{}{V}^{S}=(\prescript{W}{}{v}^{S},\prescript{W}{}{\omega}^{S})\in\mathrm{se}(2)$ (the tangent space associated with $\mathrm{SE}(2)$), where $\prescript{W}{}{v}^{S}=(d/dt)\prescript{W}{}{p}^{S}$ and $\prescript{W}{}{\omega}^{S}=(d/dt)\prescript{W}{}{\theta}^{S}$. Similarly, we represent the position of the $i$-th vertex of the slider with $\prescript{W}{}{p}^{\nu_{i}}$, and its velocity by $\prescript{W}{}{v}^{\nu_{i}}$, where $i=1,\ldots,N_{\nu}$, and $N_{\nu}$ denotes the number of slider vertices. We represent the position of the pusher in the frame of the slider by $\prescript{S}{}{p}^{P}\in\mathbb{R}^{2}$, and denote its velocity by $\prescript{S}{}{v}^{P}=(d/dt)\prescript{S}{}{p}^{P}$. We do not assign an orientation to the pusher, as it is assumed to be circular and hence invariant to rotation. When there is contact between the slider and the pusher, we will denote the position of the contact point in the slider frame by $\prescript{S}{}{p}^{c}\in\mathbb{R}^{2}$ and the velocity by $\prescript{S}{}{v}^{c}=(d/dt)\prescript{S}{}{p}^{c}$ Going forward we will use these definitions, but will often drop the frames to keep notation light. We use $(p^{S},\theta)$ and $V=(v^{S},\omega^{S})$ to denote the slider pose and spatial velocity in the world frame, and $p^{P}$, $v^{P}$, $p^{c}$, and $v^{c}$ to denote the pusher position, translational pusher velocity, contact point position, and contact point velocity in the frame of the slider.